RC Update and Run Plan for Aug 19, 2020
RC Meeting Information
RC Daily Meetings are at 4:15 PM, remote
Wednesday Aug 17 2020.
- Reference cell during SWING (08/18) and into OWL(08/19). Production on polarized He3 at the end of OWL and beginning of DAY (08/19)
- MCC preformed an energy measurement during DAY, while Hall A was preforming a Moller Measurement
- Polarization has plateaued around 37%: Recent_NMR
- Changed polarization direction to 180 degrees: Direction
- Target swap out scheduled for Friday DAY shift. Brad is coordinating with RADCON to work out the details
- 1st pass running delayed by one week (08/31)
Pending problem(s) / Things to track
- Ion chamber trips (without beam). Was fixed during SWING -- https://logbooks.jlab.org/entry/3830236
- Constantly remind the shift crew to follow the COVID rules and regulations. I have asked each outgoing shift crew to remind the oncoming shift crew.
Recent polarization measurements
- DAY (08/19) 12:50 - 36.4%
- OWL (08/19) 01:30 - 36.9%
Current Run Plan
Normal Production running
- Make sure operational restrictions are followed.
- Make sure the raster size is correct (4--4.5mm)
- Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
- Running on cell targets: raster always ON (4--4.5mm, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
- Always take 120Hz of EDTM
- Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
- Change the IHWP at least every 8 hours when running on glass cells. If Hall A hasn't changed in 8 hours, call the RC
- Run plan
- Continue taking production data on the polarized Helium-3 target in Longitudinal (180 degrees). If not able to reach requested current, pro-rate the ABUs
When there is a Down
- If Hall will have opening for 1+ hours, let Simona know as there is work to do on the SOLID cerenkov.
- Let Simona know when the details of the Friday access are worked out with RADCON.
Notices in Effect for Shift Crew
- Follow all COVID rules and regulations as well as remind each other.
- Beam positions: 
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
- Spectrometer GUI (
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
Planned Accelerator Configuration Changes
- Hall C to 1st pass on 8/31 and back to 5th pass on 9/03
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-188.8.131.52 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920